### v3

# Relational Representation Distillation

Abstract
Knowledge distillation (KD) is an effective method for transferring knowledge from a large, well-trained teacher model to a smaller, more efficient student model. Despite its success, one of the main challenges in KD is ensuring the efficient transfer of complex knowledge while maintaining the student’s computational efficiency. Unlike previous works that applied contrastive objectives promoting explicit negative instances with little attention to the relationships between them, we introduce Relational Representation Distillation (RRD). Our approach leverages pairwise similarities to explore and reinforce the relationships between the teacher and student models. Inspired by self-supervised learning principles, it uses a relaxed contrastive loss that focuses on similarity rather than exact replication. This method aligns the output distributions of teacher samples in a large memory buffer, improving the robustness and performance of the student model without the need for strict negative instance differentiation. Our approach demonstrates superior performance on CIFAR-100 and ImageNet ILSVRC-2012, outperforming traditional KD and sometimes even outperforms the teacher network when combined with KD. It also transfers successfully to other datasets like Tiny ImageNet and STL-10. Code is available at https://github.com/giakoumoglou/distillers.
1 Introduction
Knowledge distillation (KD) is a technique that facilitates the transfer of knowledge from a larger, well-trained model (teacher) to a smaller, more efficient model (student). This is achieved by minimizing the Kullback-Leibler (KL) divergence between their outputs, allowing the student model to approximate the performance of the teacher model while maintaining lower computational complexity. This process is particularly beneficial for deployment in resource-constrained environments. A critical aspect of KD is representation learning, which enables the student model to acquire meaningful feature representations that capture the underlying data distribution. Effective representation learning in KD can significantly boost the performance of the student model across various domains, such as natural language processing, computer vision, and speech recognition [22, 46, 31]. Despite these advantages, a major challenge in KD is the efficient transfer of complex knowledge from the teacher to the student model. Ensuring that the student model captures the abstract features and nuanced information present in the teacher model, without the need for similar computational capacity, remains a significant bottleneck.
Recent advancements in KD have increasingly focused on capturing richer, more nuanced knowledge from teacher models to improve the effectiveness and robustness of student models. Relational methods emphasize the importance of maintaining structural relationships between data points by minimizing distance-wise, angle-wise, or pairwise similarity losses [41, 52]. These approaches ensure that the student model learns relational information that mirrors the teacher’s representational geometry, thereby enhancing performance on downstream tasks. On the other hand, probability-based methods and methods leveraging attention mechanisms aim to align the internal feature distributions or attention maps of the teacher and student models to ensure a more targeted knowledge transfer [42, 59]. Contrastive methods, particularly Contrastive Representation Distillation (CRD), have introduced the use of memory banks and contrastive learning objectives to better preserve the teacher’s representation space [51]. By employing contrastive losses like InfoNCE [53], CRD maximizes the mutual information between teacher and student features while using a memory buffer of negative samples to encourage separation between dissimilar representations. This framework allows the student to achieve high fidelity to the teacher’s representation. However, the explicit reliance on negative sampling can introduce constraints that may limit flexibility in certain scenarios. Additionally, some methods focus on directly aligning the prediction relations or outputs of teacher and student models using correlation-based losses or $\ell_{2}$ loss functions [23, 57]. While these methods have achieved notable success, there is a growing interest in distillation techniques that balance the need for capturing relational structures without overly rigid constraints.
Our proposed method, Relational Representation Distillation (RRD), introduces a novel approach to address these challenges by maintaining relational consistency between the teacher and student models. By leveraging a large memory buffer of teacher samples to align their output distributions, our method ensures consistent relational structures, thereby enhancing the robustness and performance of the student model.
Our contributions are threefold:
  1. 1.
    We propose a novel relational consistency framework for KD, called Relational Representation Distillation (RRD), that leverages a large memory buffer and introduces a relational consistency loss. This approach aligns the similarity distributions of teacher and student outputs, maintaining relational structures and improving the robustness and generalization of the student model.
  2. 2.
    We validate the effectiveness of RRD hthrough comprehensive testing on standard benchmarks, showcasing considerable gains in both accuracy and robustness. RRD surpasses other methods with a 23.68% relative improvement [^1] Average relative improvement is calculated as: $\frac{1}{N}\sum_{i=1}^{N}\frac{\text{Acc}^{i}{\text{RRD}}-\text{Acc}^{i}{\text{KD}}}{\text{Acc}^{i}{\text{KD}}-\text{Acc}^{i}{\text{van}}}$, where $\text{Acc}^{i}{\text{RRD}}$, $\text{Acc}^{i}{\text{KD}}$, and $\text{Acc}^{i}_{\text{van}}$ represent the accuracies of RRD, KD, and vanilla training of the $i$-th student model, respectively [51]. over conventional KD. When integrated with KD, it demonstrates a 68.66% relative improvement over standard KD, underscoring its robust potential in enhancing model performance significantly.
  3. 3.
    We provide qualitative evidence through t-SNE visualizations, demonstrating that RRD effectively preserves the spatial relationships in the embedding spaces of both student and teacher models, leading to improved consistency in feature representations and enhanced student performance.
The rest of this paper is organized as follows. Section 2 reviews related work in KD and self-supervised learning. Section 3 details our proposed methodology. Section 4 presents our experimental setup and results, and Section 5 concludes the paper.
2 Related Work
Self-supervised learning. Self-supervised learning has significantly impacted representation learning by leveraging unlabeled data. Various pretext tasks have been proposed to learn these representations such as colorizing an image [61], predicting a missing patch [43], estimating the rotation angle [16], jigsaw puzzle [39, 40]. Contrastive methods further extend this concept by pulling together an anchor and a "positive" sample in the embedding space while pushing apart the anchor from many "negative" samples [26], represented by frameworks like SimCLR [8], MoCo [18, 9, 10] and other variants [56, 25]. Different from these, ReSSL [63] introduces relational self-supervised learning, which explores the relationships between data points. These approaches have inspired various KD methods, including our proposed method, which adapts relational consistency from self-supervised learning to the KD framework.
Knowledge distillation. The study of [22] on KD introduced the idea of transferring knowledge from large, complex models to smaller, more efficient models without losing their generalization capabilities. This technique involves temperature scaling in the softmax outputs to effectively capture and convey the knowledge from the teacher model. Subsequent methods have largely focused on enhancing the KD objective by incorporating additional loss functions, often complementing the standard cross-entropy and KL-divergence losses. For example, using intermediate representations or "hints" to guide the learning process [46], and aligning the attention maps of teacher and student models [59]. Other methods preserve relational knowledge between samples [52] or align the correlation structures between teacher and student models [44]. Additionally, some techniques use variational inference to improve the knowledge transfer process [1]. Further developments include focusing on structural relationships between data points to ensure the student model learns relational information [41], and maintaining the internal dynamics of neural networks during distillation [42, 21]. Other notable techniques include network compression through factor transfer [27], optimizing networks for better transfer learning [58], promoting selectivity in the distillation process [24]. Building upon the influence of self-supervised learning, CRD [51] is the first study that combines contrastive learning with KD, which aims to maximize mutual information between the teacher and student representations. CRD utilizes a memory buffer [55], which allows it to maintain a large set of negative samples. This approach enables CRD to perform contrastive learning effectively in the KD setting, enhancing the quality of transferred knowledge. Other approaches use cross-entropy with label smoothing [64], estimate propensity scores to compensate for under-represented samples [38], or apply $\ell_{2}$ loss to align teacher and student features [57]. Some techniques preserve relations between predictions [23], employ orthogonal matrices for intra-batch feature similarity [34], or automatically assign proper target layers [4]. Researchers have introduced cross-stage connection paths [7] and direct reuse of the teacher’s classifier [3]. They have also explored differentiable meta-learning [14] and many-to-one representation matching mechanisms [30]. Studies have leveraged complementary losses to maximize correlation and mutual information [36], pairwise feature kernel comparisons [17], and second-order gradient information [65]. The field has incorporated graph-based methods [33] and contrastive intra-batch losses [6]. Some approaches derive supervision from the student’s intermediate features without additional teacher models [2], while others propose training-free frameworks for architecture search [15]. Additional work focuses on normalizing logits before applying softmax and KL divergence [50].
Our method, Relational Representation Distillation (RRD), differentiates itself from state-of-the-art methods by focusing on maintaining relational consistency and alignment between the teacher and student models. Like many knowledge distillation techniques, RRD introduces an additional loss function to the training objective [51, 6, 36, 34, 4, 23, 57]. Similar to CRD [51], RRD employs a large memory buffer of teacher samples. However, unlike CRD, which uses instance discrimination, RRD relaxes this constraint to focus on aligning the relational structures of the output distributions. This approach diverges from traditional KD methods that often rely on direct alignment of logits or intermediate features. By leveraging pairwise similarities instead of explicit negative instances, RRD not only enhances the robustness and performance of the student model but also provides a more flexible and scalable solution for KD.
3 Methodology
This section presents our methodology to improve the efficiency and accuracy of KD. Our method, Relational Representation Distillation (RRD), focuses on maintaining relational consistency between the teacher and student models by leveraging a large memory buffer of teacher samples to align their output distributions. By ensuring consistent relational structures, RRD enhances the robustness and performance of the student model. Figure 1 shows an overview of the proposed RRD method.
Figure 1: Overview of the RRD method. The student network processes input $x_{i}$ to produce embeddings $z_{i}^{S}$, while the teacher network, indicated as frozen by a snowflake, processes the same input to generate embeddings $z_{i}^{T}$. The embeddings are stored in a memory buffer to align their output distributions via softmax layers. The relational consistency between the teacher and student models is enforced by leveraging the stored embeddings to train the student network.
3.1 Preliminaries
KD involves transferring knowledge from a high-capacity teacher neural network, denoted as $f^{T}$, to a more compact student neural network, $f^{S}$. Consider $x_{i}$, with $i=1,2,\ldots,N$, as the input to these networks, typically an image. We represent the outputs at the penultimate layer (just before the final classification layer, or logits) as $z_{i}^{T}=f^{T}(x_{i})$ and $z_{i}^{S}=f^{S}(x_{i})$ for the teacher and student models, respectively. The primary objective of KD is to enable the student model to approximate the performance of the teacher model while leveraging the student’s computational efficiency. The overall distillation process can be mathematically expressed as:
$\mathcal{L}=\mathcal{L}_{\text{sup}}(y_{i},z_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(z_{i}^{T},z_{i}^{S})$ (1)
where $y_{i}$ represents the true label for the input $x_{i}$ and $\lambda$ is a hyperparameter that balances the supervised loss and the distillation loss. The loss $\mathcal{L}_{\text{sup}}$ is the alignment error between the network prediction and the annotation. For example, in the image classification task [37, 48, 45, 11], it is normally a cross-entropy loss. For object detection [29, 5], it includes bounding box regression as well. The loss $\mathcal{L}_{\text{distill}}$ is the mimic error of the student network towards a pre-trained teacher network, typically implemented as KL divergence between student and teacher outputs [22].
3.2 Relational Consistency
Our method is inspired by self-supervised learning. Traditional contrastive self-supervised learning uses instance discrimination as a pretext task and relies on ($K+1$)-softmax classification, where different instances are pushed apart, and augmented views of the same instance are expected to have identical features. Applying such properties to KD imposes overly strict constraints, where a contrastive loss encourages the representations from the teacher and student models for the same input data to be similar, while simultaneously pushing apart representations from different data inputs:
$\mathcal{L}_{\text{contrast}}(z_{i}^{T},z_{i}^{S})=-\log\frac{\exp(\phi(z_{i}^{T},z_{i}^{S})/\tau)}{\sum_{j=1}^{M}\exp(\phi(z_{i}^{S},z_{j}^{T})/\tau)}$ (2)
where $\mathcal{L}_{\text{contrast}}$ is the InfoNCE [53] loss, typically employed in self-supervised methods, $\phi$ is a similarity function, $\tau$ is a temperature parameter, and $M$ is the number of negative samples.
In this way, we do not encourage explicit negative instances (those to be pushed away) for each instance; instead, we leverage the pairwise similarities to explore their relationships. We pull the features of the student $f^{S}$ and teacher $f^{T}$. As a result, our method relaxes Equation 2, where different teacher and student outputs do not always need to be pushed away from each other; and teacher and student need to share similar but not exactly the same features.
Concretely, given an input image $x_{i}$, and the outputs $z_{i}^{T}=f^{T}(x_{i})$ and $z_{i}^{S}=f^{S}(x_{i})$ for the teacher and student models, respectively, we calculate their similarity measured by $\phi(z_{i}^{T},z_{i}^{S})$. A softmax layer can be adopted to process the calculated similarities, which then produces a relationship distribution:
$p_{i}^{T}=\frac{\exp(\phi(z_{i}^{T},z_{j}^{T})/\tau_{t})}{\sum_{k=1}^{M}\exp(\phi(z_{i}^{T},z_{k}^{T})/\tau_{t})}$ (3)
where $\tau_{t}$ is the temperature parameter for the teacher network. At the same time, we can calculate the relationship distribution for the student model as:
$p_{i}^{S}=\frac{\exp(\phi(z_{i}^{S},z_{j}^{S})/\tau_{s})}{\sum_{k=1}^{M}\exp(\phi(z_{i}^{S},z_{k}^{S})/\tau_{s})}$ (4)
where $\tau_{s}$ is a different temperature parameter for the student network. We propose to push the relational consistency between $p_{i}^{T}$ and $p_{i}^{S}$, similar to [63], by minimizing the KL divergence, which can be formulated as:
$\mathcal{L}_{\text{relational}}(z_{i}^{T},z_{i}^{S})=\textit{KL}(p_{i}^{T}\parallel p_{i}^{S})=H(p_{i}^{T},p_{i}^{S})-H(p_{i}^{T})$ (5)
where KL denotes the KL divergence between $p_{i}^{T}$ and $p_{i}^{S}$. Since $p_{i}^{T}$ will be used as a target, the gradient will be clipped here to avoid the model collapsing, thus we only minimize the cross-entropy term $H(p_{i}^{T},p_{i}^{S})$ in our implementation.
However, the quality of the target similarity distribution $p_{i}^{T}$ is crucial for reliable and stable training. To ensure this, we maintain a large memory buffer $Q$, storing the feature embeddings from teacher batches. The relational consistency between the teacher and student models is enforced by aligning the similarity distributions of their outputs using the KL divergence.
We normalize the outputs $z_{i}^{T}$ and $z_{i}^{S}$ before computing the loss, ensuring that the representations lie on a unit hypersphere. To compute $\mathcal{L}_{\text{relational}}$, we further encode $z_{i}^{T}$ and $z_{i}^{S}$ using a projection head to match the dimensions. This ensures that the representations from both models are compatible for comparison and alignment.
3.3 Final Objective
The final objective function, which includes the supervised loss and standard KL divergence, is given by:
$\mathcal{L}=\mathcal{L}_{\text{sup}}(y_{i},z_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(z_{i}^{T},z_{i}^{S})+\beta\cdot\mathcal{L}_{\text{relational}}(z_{i}^{T},z_{i}^{S})$ (6)
where $\beta$ is a hyperparameter that balances the KD loss $\mathcal{L}_{\text{relational}}$. We experiment with $\beta$ in ablation studies to understand its impact on the final performance.
4 Experiments
We evaluate our RRD framework in the KD task of model compression of a large network to a smaller one, following the same protocol as [22, 46, 59, 52, 44, 1, 41, 42, 21, 27, 58, 24, 51]. This approach is consistent with the standard practice in the field, allowing for fair comparison across methods.
Datasets. (1) CIFAR-100 [28] contains 50,000 training images with 500 images per class and 10,000 test images. (2) ImageNet ILSVRC-2012 [13] includes 1.2 million images from 1,000 classes for training and 50,000 for validation. (3) STL-10 [12] consists of a training set of 5,000 labeled images from 10 classes and 100,000 unlabeled images, and a test set of 8,000 images. (4) Tiny ImageNet (TIN-200) [13] has 200 classes, each with 500 training images and 50 validaton images.
Setup. We experiment on CIFAR-100 and ImageNet with student-teacher combinations of various capacity, such as ResNet [19] or Wide ResNet (WRN) [60], VGG [49], MobileNet [47], and ShuffleNet [62, 32] (more details about the network architectures are described in the supplementary material). We use $M=16384$ negative samples as in [51] (we ablate $M$ in Section 4.6) and set the temperature parameter of the student to $\tau_{s}=0.04$ and of the teacher to $\tau_{t}=0.07$, also used in [63] that employs dual temperature (we also ablate both temperatures in Section 4.6). Following [51], both the student and teacher outputs are projected to a 128-dimensional space. We use a projection head of a single linear layer, followed by $\ell_{2}$ normalization. We train for a total of 240 epochs. The hyperparameter $\lambda$ is set to $1.0$ for the KL divergence loss to maintain consistency with the baseline CRD method [51], facilitating fair comparison. This value of $\lambda$ is also fixed to $1.0$ in other methods such as [3, 4, 51]. The loss coefficient $\beta$ is set to $1.0$ (we study the impact of $\beta$ in Section 4.6). More details on the training details are described in the supplementary material.
Comparison. We compare our approach to the following state-of-the-art methods: (1) KD [22]; (2) FitNets [46]; (3) AT [59]; (4) SP [52]; (5) CC [44]; (6) VID [1]; (7) RKD [41]; (8) PKT [42]; (9) AB [21]; (10) FT [27]; (11) FSP [58]; (12) NST [24]; (13) CRD [51]; (14) OFD [20]; (15) WSLD [64]; (16) IPWD [38]. In the supplementary material, we include additional methods: (17) SRRL [57]; (18) SemCKD [4]; (19) ReviewKD [7]; (20) SimKD [3]; (21) DistPro [14]; (22) NORM [30]; (23) ITRD [36]; (24) FKD [17]; (25) CRCD [65]; (26) EGA [33]; (27) WCoRD [6].
Role of the projection layer. Following [35], we use a projection head consisting of a single linear layer, followed by $\ell_{2}$ normalization. The projection layer is not merely a dimension-matching tool but plays a crucial role in KD. It effectively transfers knowledge by implicitly encoding relational information from previous samples [35]. The weights within this projection layer capture crucial relational data, which is fundamental for conveying information about the correlation between features, thus significantly influencing the training dynamics.
4.1 Results on CIFAR-100
Table 1 and Table 2 present the top-1 accuracies of student networks trained using different distillation techniques across various teacher-student architectural pairings. Table 1 examines pairings where both student and teacher models share similar architectural styles, while Table 2 focuses on cross-architecture distillations. Our proposed loss consistently outperforms the conventional KD technique, as indicated by the green upward arrows ($\uparrow$). While the standalone performance of our method is comparable to CRD, its integration with KD not only achieves higher accuracies but in some cases, surpasses the performance of the teacher networks, such as in the distillation of WRN-40-2 to ShuffleNet-v1. The enhanced performance of our distillation method can be credited to multiple factors that collectively improve the transfer of knowledge from the teacher to the student model. Our approach uses a unique loss function that complements KD’s primary focus on matching the softened output logits of the teacher and student. We introduce an additional layer of representational alignment that ensures not only the final outputs but also the intermediate feature representations of the student closely match those of the teacher. This dual focus allows the student model to mimic the teacher’s outputs and develop more robust and generalizable internal representations.
Teacher WRN-40-2 WRN-40-2 resnet-56 resnet-110 resnet-110 resnet-32x4 VGG-13
Student WRN-16-2 WRN-40-1 resnet-20 resnet-20 resnet-32 resnet-8x4 VGG-8
Teacher 75.61 75.61 72.34 74.31 74.31 79.42 74.64
Student 73.26 71.98 69.06 69.06 71.14 72.50 70.36
KD [22] 74.92 73.54 70.66 70.67 73.08 73.33 72.98
FitNet [46] 73.58 ($\downarrow$) 72.24 ($\downarrow$) 69.21 ($\downarrow$) 68.99 ($\downarrow$) 71.06 ($\downarrow$) 73.50 ($\downarrow$) 71.02 ($\downarrow$)
AT [59] 74.08 ($\downarrow$) 72.77 ($\downarrow$) 70.55 ($\downarrow$) 70.22 ($\downarrow$) 72.31 ($\downarrow$) 73.44 ($\downarrow$) 71.43 ($\downarrow$)
SP [52] 73.83 ($\downarrow$) 72.43 ($\downarrow$) 69.67 ($\downarrow$) 70.04 ($\downarrow$) 72.69 ($\downarrow$) 72.94 ($\downarrow$) 72.68 ($\downarrow$)
CC [44] 73.56 ($\downarrow$) 72.21 ($\downarrow$) 69.63 ($\downarrow$) 69.48 ($\downarrow$) 71.48 ($\downarrow$) 72.97 ($\downarrow$) 70.81 ($\downarrow$)
VID [1] 74.11 ($\downarrow$) 73.30 ($\downarrow$) 70.38 ($\downarrow$) 70.16 ($\downarrow$) 72.61 ($\downarrow$) 73.09 ($\downarrow$) 71.23 ($\downarrow$)
RKD [41] 73.35 ($\downarrow$) 72.22 ($\downarrow$) 69.61 ($\downarrow$) 69.25 ($\downarrow$) 71.82 ($\downarrow$) 71.90 ($\downarrow$) 71.48 ($\downarrow$)
PKT [42] 74.54 ($\downarrow$) 73.45 ($\downarrow$) 70.34 ($\downarrow$) 70.25 ($\downarrow$) 72.61 ($\downarrow$) 73.64 ($\uparrow$) 72.88 ($\downarrow$)
AB [21] 72.50 ($\downarrow$) 72.38 ($\downarrow$) 69.47 ($\downarrow$) 69.53 ($\downarrow$) 70.98 ($\downarrow$) 73.17 ($\downarrow$) 70.94 ($\downarrow$)
FT [27] 73.25 ($\downarrow$) 71.59 ($\downarrow$) 69.84 ($\downarrow$) 70.22 ($\downarrow$) 72.37 ($\downarrow$) 72.86 ($\downarrow$) 70.58 ($\downarrow$)
FSP [58] 72.91 ($\downarrow$) n/a 69.95 ($\downarrow$) 70.11 ($\downarrow$) 71.89 ($\downarrow$) 72.62 ($\downarrow$) 70.33 ($\downarrow$)
NST [24] 73.68 ($\downarrow$) 72.24 ($\downarrow$) 69.60 ($\downarrow$) 69.53 ($\downarrow$) 71.96 ($\downarrow$) 73.30 ($\downarrow$) 71.53 ($\downarrow$)
CRD [51] 75.48 ($\uparrow$) 74.14 ($\uparrow$) 71.16 ($\uparrow$) 71.46 ($\uparrow$) 73.48 ($\uparrow$) 75.51 ($\uparrow$) 73.94 ($\uparrow$)
CRD+KD [51] 75.64 ($\uparrow$) 74.38 ($\uparrow$) 71.63 ($\uparrow$) 71.56 ($\uparrow$) 73.75 ($\uparrow$) 75.46 ($\uparrow$) 74.29 ($\uparrow$)
OFD [20] 75.24 ($\uparrow$) 74.33 ($\uparrow$) 70.38 ($\uparrow$) n/a 73.23 ($\uparrow$) 74.95 ($\uparrow$) 73.95 ($\uparrow$)
WSLD [64] n/a 73.74 ($\uparrow$) 71.53 ($\uparrow$) n/a 73.36 ($\uparrow$) 74.79 ($\uparrow$) n/a
IPWD [38] n/a 74.64 ($\uparrow$) 71.32 ($\uparrow$) n/a 73.91 ($\uparrow$) 76.03 ($\uparrow$) n/a
RRD (ours) 75.33 ($\uparrow$) 73.55 ($\uparrow$) 70.71 ($\uparrow$) 70.72 ($\uparrow$) 73.10 ($\uparrow$) 74.48 ($\uparrow$) 73.99 ($\uparrow$)
RRD+KD (ours) 75.66 ($\uparrow$) 74.67 ($\uparrow$) 72.19 ($\uparrow$) 71.74 ($\uparrow$) 73.54 ($\uparrow$) 75.08 ($\uparrow$) 74.32 ($\uparrow$)
Table 1: Test top-1 accuracy (%) of student networks on CIFAR-100, comparing students and teachers of the same architecture using various distillation methods. $\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance. The values in bold indicate the maximum of each column.
Teacher VGG-13 ResNet-50 ResNet-50 ResNet-32x4 ResNet-32x4 WRN-40-2
Student MobileNet-v2 MobileNet-v2 VGG-8 ShuffleNet-v1 ShuffleNet-v2 ShuffleNet-v1
Teacher 74.64 79.34 79.34 79.42 79.42 75.61
Student 64.6 64.6 70.36 70.5 71.82 70.5
KD [22] 67.37 67.35 73.81 74.07 74.45 74.83
FitNet [46] 64.14 ($\downarrow$) 63.16 ($\downarrow$) 70.69 ($\downarrow$) 73.59 ($\downarrow$) 73.54 ($\downarrow$) 73.73 ($\downarrow$)
AT [59] 59.40 ($\downarrow$) 58.58 ($\downarrow$) 71.84 ($\downarrow$) 71.73 ($\downarrow$) 72.73 ($\downarrow$) 73.32 ($\downarrow$)
SP [52] 66.30 ($\downarrow$) 68.08 ($\downarrow$) 73.34 ($\downarrow$) 73.48 ($\downarrow$) 74.56 ($\uparrow$) 74.52 ($\downarrow$)
CC [44] 64.86 ($\downarrow$) 65.43 ($\downarrow$) 70.25 ($\downarrow$) 71.14 ($\downarrow$) 71.29 ($\downarrow$) 71.38 ($\downarrow$)
VID [1] 65.56 ($\downarrow$) 67.57 ($\downarrow$) 70.30 ($\downarrow$) 73.38 ($\downarrow$) 73.40 ($\downarrow$) 73.61 ($\downarrow$)
RKD [41] 64.52 ($\downarrow$) 64.43 ($\downarrow$) 71.50 ($\downarrow$) 72.28 ($\downarrow$) 73.21 ($\downarrow$) 72.21 ($\downarrow$)
PKT [42] 67.13 ($\downarrow$) 66.52 ($\downarrow$) 73.01 ($\downarrow$) 74.10 ($\uparrow$) 74.69 ($\uparrow$) 73.89 ($\downarrow$)
AB [21] 66.06 ($\downarrow$) 67.20 ($\downarrow$) 70.65 ($\downarrow$) 73.55 ($\downarrow$) 74.31 ($\downarrow$) 73.34 ($\downarrow$)
FT [27] 61.78 ($\downarrow$) 60.99 ($\downarrow$) 70.29 ($\downarrow$) 71.75 ($\downarrow$) 72.50 ($\downarrow$) 72.03 ($\downarrow$)
NST [24] 58.16 ($\downarrow$) 64.96 ($\downarrow$) 71.28 ($\downarrow$) 74.12 ($\uparrow$) 74.68 ($\uparrow$) 76.09 ($\uparrow$)
CRD [51] 69.73 ($\uparrow$) 69.11 ($\uparrow$) 74.30 ($\uparrow$) 75.11 ($\uparrow$) 75.65 ($\uparrow$) 76.05 ($\uparrow$)
CRD+KD [51] 69.94 ($\uparrow$) 69.54 ($\uparrow$) 74.58 ($\uparrow$) 75.12 ($\uparrow$) 76.05 ($\uparrow$) 76.27 ($\uparrow$)
OFD [20] 69.48 ($\uparrow$) 69.04 ($\uparrow$) n/a 75.98 ($\uparrow$) 76.82 ($\uparrow$) 75.85 ($\uparrow$)
WSLD [64] n/a 68.79 ($\uparrow$) 73.80 ($\uparrow$) 75.09 ($\uparrow$) n/a 75.23 ($\uparrow$)
RRD (ours) 67.93 ($\uparrow$) 68.84 ($\uparrow$) 74.01 ($\uparrow$) 74.11 ($\uparrow$) 74.80 ($\uparrow$) 74.98 ($\uparrow$)
RRD+KD (ours) 69.98 ($\uparrow$) 69.13 ($\uparrow$) 74.26 ($\uparrow$) 75.18 ($\uparrow$) 76.83 ($\uparrow$) 76.31 ($\uparrow$)
Table 2: Test top-1 accuracy (%) of student networks on CIFAR-100 involving students and teachers from different architectures, using various distillation methods. $\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance. The values in bold indicate the maximum of each column.
4.2 Results on ImageNet
Table 3 presents the top-1 accuracies of student networks trained with various distillation techniques across different teacher-student architectural pairings. These findings affirm the scalability of our RRD method on large datasets like ImageNet, highlighting its ability to effectively distill complex models. Our approach achieves competitive results, surpassing KD across all tested architectures. Furthermore, RRD shows improvement across different architectures, demonstrating its effectiveness in various distillation scenarios. The combination of RRD with KD further improves results among the compared techniques in most cases.
Teacher ResNet-34 ResNet-50 ResNet-50
Student ResNet-18 ResNet-18 MobileNet
Teacher 73.31 76.16 76.16
Student 69.75 69.75 69.63
KD [22] 70.67 71.29 70.49
AT [59] 71.03 ($\uparrow$) 71.18 ($\downarrow$) 70.18 ($\downarrow$)
SP [52] 70.62 ($\uparrow$) 71.08 ($\downarrow$) n/a
CC [44] 69.96 ($\downarrow$) n/a n/a
VID [1] n/a 71.11 ($\downarrow$) n/a
RKD [41] 70.40 ($\downarrow$) n/a 68.50 ($\downarrow$)
AB [21] n/a n/a 68.89 ($\downarrow$)
FT [27] n/a n/a 69.88 ($\downarrow$)
FSP [58] 70.58 ($\downarrow$) n/a n/a
NST [24] 70.29 ($\downarrow$) n/a n/a
CRD [51] 71.17 ($\uparrow$) 71.25 ($\downarrow$) 69.07 ($\downarrow$)
OFD [20] 71.03 ($\uparrow$) n/a 71.33 ($\uparrow$)
WSLD [64] 72.04 ($\uparrow$) n/a 71.52 ($\uparrow$)
IPWD [38] 71.88 ($\uparrow$) n/a 72.65 ($\uparrow$)
RRD (ours) 71.22 ($\uparrow$) 71.33 ($\uparrow$) 70.66 ($\uparrow$)
RRD+KD (ours) 71.40 ($\uparrow$) 71.51 ($\uparrow$) 71.83 ($\uparrow$)
Table 3: Test top-1 (%) on ImageNet validation test using various distillation methods. The table compares students and teachers of the same and different architecture. $\uparrow$ denotes outperformance over KD and $\downarrow$ denotes underperformance. The values in bold indicate the maximum of each column while underlined values mark the second best.
4.3 Capturing Inter-class Correlations
Cross-entropy loss overlooks the relationships among class logits in a teacher network, often resulting in less effective knowledge transfer. Distillation techniques that use "soft targets", such as those described by [22], have successfully captured these relationships, improving student model performance. Figure 2 assesses the effectiveness of different distillation methods on the CIFAR-100 KD task using WRN-40-2 as the teacher and WRN-40-1 as the student. We compare students trained without distillation, with attention transfer [59], with KL divergence [22], and with our proposed RRD method. Our findings show that RRD achieves close alignment between teacher and student logits, as evidenced by reduced differences in their correlation matrices. While RRD does not match CRD [51] in terms of exact correlation alignment, it significantly enhances learning efficiency and reduces error rates. The smaller discrepancies between teacher and student logits indicate that the RRD objective captures a substantial portion of the correlation structure in the logits, resulting in lower error rates, though CRD achieves a slightly closer match.
(a) Student: vanilla
(b) Student: AT
(c) Student: KD
(d) Student: RRD (ours)
Figure 2: Comparison of correlation matrix differences between teacher and student logits across various distillation methods on the CIFAR-100 task. Subfigures show results for (a) students trained without distillation, (b) with attention transfer [59], (c) with KL divergence [22], and (d) with our RRD method, highlighting better matching between student’s and teacher’s correlations. Results have been re-implemented according to [51].
4.4 Transferability of Representations
Our study investigates how knowledge is transferred from a larger teacher network (WRN-40-2) to a smaller student network (WRN-16-2), aiming to develop versatile representations suitable for a range of tasks and datasets. We apply this technique by having the student network either learn directly from the CIFAR-100 dataset or via distillation. In this setup, the student network is employed as a static feature extractor for images from the STL-10 and TIN-200 datasets, both adjusted to a $32\times 32$ resolution. We evaluate the adaptability of these features by training a linear classifier on the final feature layer to conduct classifications with 10 categories for STL-10 and 200 categories for TIN-200. We document the impact of various distillation approaches on the transferability of these features in Table 4. Our findings reveal that, except for the FitNet method, all distillation techniques significantly improve the feature transferability on both datasets. Notably, while the teacher network achieves the highest performance on CIFAR-100, its features show the least effective transfer, likely due to training data specificity. Conversely, the student network using a combination of RRD and KD distillation techniques not only equates to the teacher’s CIFAR-100 performance but also surpasses it in transferability, showing improvements of 3.1% and 3.1% in STL-10 and TIN-200 respectively.
Teacher Student KD AT FitNet CRD CRD+KD RRD RRD+KD
[…]
Table 4: Test top-1 accuracy (%) of WRN-16-2 (student) distilled from WRN-40-2 (teacher). In this setup, the representations learned from the CIFAR-100 dataset are transferred to the STL-10 and TIN-200 datasets. The network is frozen, and a linear classifier is trained on the last feature layer to perform classification with 10 classes (STL-10) or 200 classes (TIN-200). Results adapted from [51]. The values in bold indicate the maximum of each row.
4.5 Visualization of t-SNE Embeddings
We provide t-SNE [54] visualizations to compare the embeddings generated by various KD methods and the teacher network on the CIFAR-100 dataset. Figure 3 displays the embeddings from the teacher network, a WRN-40-2, and the student network, WRN-40-1, under standard training as well as distillation using AT and RRD where we limit the dataset to the first 10 classes of CIFAR-100 to offer a clearer understanding of the embedding space. We observe improved consistency in the embedding distributions between the teacher and student networks, indicating that RRD effectively transfers the knowledge of the teacher’s feature space to the student. Our relational consistency approach ensures that the spatial relationships in the embedding spaces of both the student and teacher models are preserved. This alignment not only enhances the student’s performance but also maintains the integrity of the feature representations learned by the teacher.
(a) Teacher
(b) Student: vanilla
(c) Student: AT
(d) Student: RRD (ours)
Figure 3: t-SNE visualizations of embeddings from the teacher network and student networks trained using different distillation techniques on the first 10 classes of the CIFAR-100 dataset.
4.6 Ablation Study
(a) Effect of varying $M$
(b) Effect of varying $\tau_{s}$ and $\tau_{t}$
(c) Effect of varying loss coefficient $\beta$
Figure 4: Ablation study results on CIFAR-100 using WRN-40-2 as the teacher and WRN-16-2 as the student. (a) Effect of the number of negatives $M$ on performance. Increasing $M$ generally improves performance, but the difference becomes negligible beyond $M=4096$. (b) Effect of temperature parameters $\tau_{s}$ and $\tau_{t}$. Lower $\tau_{s}$ than $\tau_{t}$ results in better performance, with extremely high or low temperatures leading to sub-optimal solutions. (c) Effect of loss coefficient $\beta$ on performance (logarithmic scale). Optimal $\beta$ values range between 1 and 2.
There are two main hyperparameters in our objective: the number of negative samples $M$ in the memory buffer $\mathcal{Q}$ of Equation 3 and Equation 4; and temperature parameters $\tau_{s}$ and $\tau_{t}$ of Equation 3 and Equation 4 that modulate the softmax probability. We also ablate the hyperparameter $\beta$ that balances the KD loss. For the ablation study, we adopt WRN-40-2 as the teacher and WRN-16-2 as the student. Experiments are conducted on CIFAR-100, and the results are shown in Figure 4.
Number of negatives $\bm{M}$. We validated different values for $M$: 256, 1024, 2048, 4096, 8192, 16384, and 32768. As shown in Figure 4(a), increasing $M$ leads to improved performance. However, the difference in error rate between $M=4096$ and $M=16384$ is less 0.5%. Therefore, we use $M=16384$ for reporting the accuracy, while in practice lower $M$ should suffice. Going beyond $M=16384$ proves to harms performance.
Temperature $\bm{\tau_{s}}$ and $\bm{\tau_{t}}$. We varied $\tau_{s}$ and $\tau_{t}$, considering all the permutations of the numbers 0.04, 0.07, and 0.2 taken two at a time, considering that order matters. As Figure 4(b) illustrates, both extremely high or low temperatures lead to a sub-optimal solution. Also, as expected, a lower $\tau_{s}$ than $\tau_{t}$ leads to better performance. This improvement is attributed to the sharper predictions from the student model with a lower temperature, resulting in more confident and distinct class probabilities. This sharpness helps the student model align better with the teacher’s guidance, enhance generalization, provide a more informative learning signal, and reduce overfitting. The sweep spot lies at $\tau_{s}=0.04$ and $\tau_{t}=0.07$ (red area of Figure 4(b)).
Loss coefficient $\bm{\beta}$. We varied $\beta$ from 0.1 to 100. As Figure 4(c) illustrates, both extremely high or low $\beta$ lead to a sub-optimal solution. In general, $\beta$ between 1 and 2 works well on CIFAR-100.
Computational cost. The RRD loss method introduces an additional 5.24 MFLOPs to the computational workload of a ResNet-50 model. This represents approximately 0.262% of the original 2 GFLOPs computational baseline. Despite the added computational complexity, the impact on training time is minimal, allowing us to maintain efficiency in practical training scenarios such as on CIFAR-100. The memory bank for RRD stores all 128-dimensional features of up to 16,384 CIFAR-100 images, consuming about 8 MB of GPU memory.
5 Conclusions
Our method offers a significant advancement in KD by maintaining relational consistency between teacher and student models. RRD leverages a large memory buffer of teacher samples to align their output distributions, ensuring consistent relational structures throughout the learning process. Unlike traditional approaches, RRD uses pairwise similarities and a relaxed contrastive loss to enhance the robustness and performance of the student model without explicit negative instances. Through comprehensive testing on CIFAR-100, ImageNet, TIN-200, and STL-10 datasets, RRD consistently outperforms traditional KD techniques and state-of-the-art methods, demonstrating substantial improvements in accuracy, robustness, and transferability of learned representations.